# Large Language Models as Proxies for Theories of Human Linguistic Cognition

### v1

@@
Abstract
@@
1 Introduction
The question of whether current large language models (LLMs) might be relevant scientifically for the study of human linguistic cognition (HLC) is often framed in terms of a perceived threat that LLMs pose to generative linguistics. On the one hand, some researchers have suggested that these models discredit prevailing views within linguistics (Piantadosi, 2023), while others have suggested that LLMs can teach us nothing about HLC (Chomsky et al., 2023, Moro et al., 2023a).
Here we chart a more collaborative path forward for the integration of LLMs into the study of HLC. While the idea that current LLMs are themselves good theories of HLC is untenable (as reviewed below), we note that LLMs are a potentially useful tool that can inform the evaluation of theories of HLC. In this capacity, LLMs can serve as proxies for explicit theories of HLC that are different from LLMs in important ways but are still relatively unbiased with respect to linguistic representations and learning. We take it that much of the recent literature on the potential cognitive relevance of LLMs takes current models to support relatively linguistically neutral theories of this kind. To the extent that this use of LLMs as proxies is justified, they can assist in comparing such relatively unbiased theories of HLC with more traditional theories of HLC coming out of theoretical linguistics, where the combination of representations and learning is often highly biased in favor of linguistic patterns.
We start in section 2 by briefly reviewing some considerations pertaining to the role of LLMs in studying HLC. We start with a reminder that the study of HLC, like the rest of empirical science, proceeds through inference to the best explanation, with explicit theories competing in light of known observations. We review the recent idea that LLMs themselves can serve as better explanations than generative theories and then summarize some of the reasons noted in the literature to reject this idea. We then consider the idea mentioned above that the best explanation is that HLC is quite different from current LLMs but still involves representations and learning that are relatively linguistically neutral. It is the potential of LLMs to serve as proxies for such a relatively linguistically-neutral theory that will preoccupy us in the remainder of the paper, and in particular the question of whether evidence from LLMs can bear on the question of whether the linguistically-neutral theory is a better explanation than generative approaches. In section 3 we ask – building on Wilcox et al. 2023 and Lan et al. 2024b, among others – whether evidence from LLMs suggests that a relatively linguistically-neutral theory of HLC might succeed in acquiring various linguistic patterns from a developmentally-realistic corpus. In section 4 we ask – building on Kallini et al. 2024 and others – whether evidence from LLMs suggests that a relatively linguistically-neutral theory of HLC makes attested languages easier to acquire than various unattested variants of those languages. In both cases, we will see that the evidence from LLMs gives no reason to think that a linguistically-neutral theory of HLC could account for the data, which in turn means that current LLMs contribute little to the ability of such linguistically-neutral theories to challenge generative theories.
@@ 2 Background: LLMs, HLC, and the inference to the best explanation
Being part of empirical science, the study of HLC proceeds through inference to the best explanation: explicit theories are evaluated in light of the observations, and those theories that explain the data best are preferred to the other contenders. The theories coming out of generative linguistics over the past 70 years or so are routinely evaluated in this way. These theories are often strongly linguistically biased, in the sense that they make it relatively easy to represent and learn knowledge of the kind found across languages and hard or impossible to represent or learn many other patterns. Let us use $H_{1}$ to refer to some concrete instantiation of the generative approach. We will return to the biases embodied by the architecture and representations of $H_{1}$ in the following sections.
More recently, Piantadosi (2023) has suggested that current LLMs should be treated as explicit theories of HLC that should be taken seriously as part of this evaluation. Following Fox and Katzir (2024) we refer to this new approach to HLC as the LLM Theory, and we discuss it in Section 2.1, through a concrete instantiation we will call $H_{2}$. Perhaps in light of the glaring inadequacies of the LLM Theory, some recent literature, including Wilcox et al. 2023 and Mahowald et al. 2024, has advocated for a revision of generative linguistics on the basis of LLMs but without committing to the LLM Theory. We refer to this idea as the Proxy View:
@@
Since the literature that promotes the Proxy View generally does not make its supported theory explicit, our ability to evaluate it at present is limited. For concreteness, we will outline what such a target theory might look like, a theory that we will refer to as $H_{3}$ and that addresses at least some of the most immediate inadequacies of the LLM Theory but is still more linguistically neutral than generative proposals. But we can of course have no certainty that this is what Wilcox et al. (2023), Mahowald et al. (2024), and others have in mind. We will therefore focus on methodological considerations that pertain to the evaluation of the target theory once it is provided and show how LLMs might bear on this evaluation. We show that, at least at present, LLMs lend no support to the Proxy View.
2.1 The LLM Theory and its inadequacy
As mentioned, Piantadosi (2023) argues for the LLM Theory of HLC. That is, he suggests that current LLMs, taken as real theories of HLC, are a better explanation of the known observations than anything coming out of the generative tradition. This claim is incorrect, for reasons discussed in detail in the literature and rehearsed briefly here. Following Katzir 2023 and Fox and Katzir 2024 we structure our brief discussion of the failure of the LLM Theory around three key issues: (a) competence vs. performance, (b) correctness vs. probability, and (c) learning and representations. [^1] See Kodner et al. 2023, Moro et al. 2023b, Rawski and Baumont 2023, and others for further critical discussion of the LLM Theory.
Competence vs. performance. Humans sometimes struggle with sentences that, given more time and focus, they ultimately accept. Familiar examples include center embedding, as in "The mouse that the cat that the dog chased bit died". Humans also sometimes initially accept sentences that, upon further thought, they reject. Familiar examples include variants of sentences such as the above but in which the verbs and the nouns are not balanced, as in "The mouse that the cat that the dog chased died". Other examples include so-called agreement attraction, as in "The keys to the cabinet is on the table".
The best explanation to these and other examples, going back to Yngve 1960 and Miller and Chomsky 1963, distinguishes linguistic competence from performance. The competence of English speakers, for example, licenses center-embedded structures regardless of depth, and it requires agreement between verbs and subjects regardless of intervening nouns. Performance factors get in the way and make center embedding and agreement across intervening nouns difficult and error prone. Among other things, the explanation in terms of a distinction between competence and performance accounts for the way center embedding can become easier or harder given various manipulations (tiredness, noise, parallel tasks, etc.).
Differently from common generative proposals, the design and behavior of current LLMs does not suggest a meaningful distinction between competence and performance of the kind discovered in humans: there is no reason to think that more (or less) time and working memory would make them change their behavior on center embedding or agreement. This leaves current LLMs with no good explanation for the increasing difficulty of sentences with increasing levels of center embedding or for the other observations that have been taken to support the distinction between competence and performance.
Correctness vs. likelihood. Humans recognize some sentences as correct but unlikely (one example again is center embedding as in the example above), and they recognize some other sentences as incorrect but likely (again, unbalanced variants of center embedding and agreement attraction, among other examples). LLMs provide straightforward information regarding their representation of likelihood. But it remains unclear whether they even have a distinct representation of correctness. [^2] A potentially important area of research concerns understanding the inner workings of LLMs, sometimes relating the discovered dynamics to linguistic patterns. See Lakretz et al. 2019 and Ravfogel et al. 2021, among others. It is conceivable that such investigations will eventually uncover a notion of correctness within LLMs.
[…]
@@ 2.1 The LLM Theory and its inadequacy
Generative theories of HLC typically derive this inductive leap by assuming that children are born with a combination of representations and learning biases that favor the representation of constituency-based dependencies over dependencies that rely on linear order (in some approaches by preventing the latter from being represented in the first place). Current LLMs, on the other hand, seem to provide no basis for this inductive leap. In fact, recent work by Lan et al. (2022, 2024a) casts doubt on the ability of networks trained by current methods to even acquire the basic notion of constituency, let alone to explain the preference for constituency-based generalizations over linear-based ones. Not surprisingly, then, Yedetore et al. (2023) find evidence suggesting that at least some neural networks fail to acquire the correct constituency-based generalization from their training data. We will return to APS in section 3, where we will consider the possible role of LLMs in reasoning about the inductive leaps predicted by non-LLM theories of HLC.
@@
Evidence about the representations and learning biases of humans also comes from the highly skewed distribution of linguistic patterns cross-linguistically. We can again consider constituency (following Fox and Katzir 2024), though we emphasize that a very wide range of cross-linguistic patterns have been discussed in the linguistic literature. We just mentioned that constituency is part of the best explanation of the competence of English speakers, noting its role in the formation of yes-no questions. Constituency is central to the explanation of many other aspects of English, including center-embedding and agreement mentioned earlier, wh-movement discussed in section 3 below, the computation of meaning, and more. But it seems clear that this is not just an accident of actual English: constituency is part of the best explanation of almost all languages that have been studied closely, which suggests that humans are born with representations and learning biases that favor constituency. Again, the representations and learning biases of current LLMs might not even be able to support the acquisition of constituency, let alone explain its cross-linguistic prevalence. We will return to cross-linguistic evidence in section 4, where we will consider the possible role of LLMs in reasoning about the typological patterns predicted by non-LLMs theories of HLC.
@@ 2.2 LLMs as proxies for good learners
But if the LLM Theory is discredicted, does this mean that LLMs are irrelevant to the scientific study of HLC? As mentioned above, the Proxy View explains why the answer is no: in principle LLMs can serve as proxies for theories of a third kind, theories that involve relatively linguistically-neutral representations and learning (and are different in this regard from generative theories) but that are not vulnerable to the most immediate arguments against the LLM Theory.
Here we consider two such uses of LLMs as proxies. Following Lan et al. 2024b we use LLMs as tools for reasoning about whether a relatively linguistically-neutral theory of HLC predicts a given inductive leap based on a given amount of linguistic exposure. And following Kallini et al. 2024 we use LLMs as tools for reasoning about whether a relatively linguistically-neutral theory of HLC derives a learning asymmetry that might be part of an explanation of a typological asymmetry.
@@
At this point we should mention again an immediate obstacle to the proxy use of LLMs: while recent literature, including Wilcox et al. (2023) and Mahowald et al. (2024), suggests that the performance of LLMs supports linguistically-neutral theories of HLC, this literature does not provide much by way of specifics. Until such specifics are provided, expressions of hope about the undisclosed theories remain just that.
Still, we can outline some methodological considerations regarding the possible role of current LLMs in evaluating a linguistically-neutral theory of HLC once such a theory is made sufficiently explicit.
Let us illustrate with a concrete target theory, though we emphasize that this is for presentational purposes only and that we obviously cannot tell whether this target theory is at all close to what the literature supporting the Proxy View envisions. Consider then the following theory of HLC, which we will call $H_{3}$. According to $H_{3}$, the syntax is written and understood as Multiple Context-Free Grammar: any MCFG can be written, using the formalism of Seki et al. (1991), and no other grammar can be written. The semantics in $H_{3}$ is model-theoretic and is computed compositionally based on the syntax, while the syntax is blind to most of the semantics. Learning according to this theory follows the principle of Minimum Description Length (MDL; Rissanen, 1978). And processing relies on a parser with access to very little working memory and that can output not just parses (for grammatical inputs) but also probabilities.
$H_{3}$ is quite different from the LLM Theory and does not share the inadequacies of that theory reviewed above. For one thing, $H_{3}$ has a meaningful distinction between competence and performance: by relying on a limited amount of working memory, the parser will run into trouble with some of the same grammatically correct sentences that humans struggle with such as those with center embedding. $H_{3}$ also has a meaningful distinction between correct and probable. As to representations and learning, the combination of (a) MCFG as the representational framework for syntax and (b) MDL as the learning criterion will allow the learner to represent and learn constituency and perhaps also constituency-based dependencies. And the combination of those representations with compositional semantics can go a long way toward explaining the centrality of constituency across languages. As to modularity, this is baked into the assumptions mentioned earlier about the information available to the syntax. But while $H_{3}$ is very different from the LLM Theory, it is also more linguistically-neutral than common generative theories of HLC such as $H_{1}$. The representational system of $H_{1}$ would mostly likely include a specific preparation for grammars that involve linguistic patterns attested across langauges, such as constraints on movement commonly known as syntactic islands (Ross 1967), and including the phenomena we discuss in the following section. While $H_{3}$ makes it possible to state such grammars, it does not postulate anything specific to these patterns inside the theory, and does not provide further explicit biases beyond those present in a constituency-based syntax and model-theoretic semantics. It is therefore conceivable that the performance of LLMs might be informative to some extent about what can be learned and under what conditions under $H_{3}$, though adequately justifying the Proxy View for $H_{3}$ would require careful argumentation that goes beyond the present outline.
@@
Explanation for learning assymetries $\checkmark$ $\times$ $\checkmark$
@@
Table 1: Comparison of the three theories outlined above on key issues, following Fox and Katzir 2024.
As mentioned, we have no way of telling whether $H_{3}$ is the theory of HLC that works such as Wilcox et al. 2023 and Mahowald et al. 2024 seem to have tried to promote on the basis of LLM performance. None of our methodological discussion below depends in any way on the specific properties of $H_{3}$: this theory can serve as a concrete example of a relatively linguistically-neutral theory of HLC if having such an example is helpful, but the reader should feel perfectly free to ignore it. It should also be clear that any serious attempt to promote $H_{3}$ or any other theory of HLC - as the best explanation would need to examine it in light of the whole body of work within theoretical linguists since the 1950s and that linguists have relied on to argue for theories of HLC that incorporate strong linguistic biases. Since our goal here is merely to illustrate the potential use of LLMs as proxies for linguistically-neutral theories, we will proceed as if this task has already been carried out.

### v2

@@
Large Language Models as Proxies for Theories of Human Linguistic Cognitionthanks: We thank Danny Fox, Fred Landman, Roger Levy, Eyal Marco, and audiences at Leibniz ZAS, MIT, and TAU, as well as the reviewers and editors at JLM. RK has been supported by ISF grant #1083/23 and by the Alexander von Humboldt Foundation. This project was provided with computer and storage resources by GENCI at IDRIS thanks to the grant AD011013783R2 on the supercomputer Jean Zay’s V100 and CSL partitions.
Imry Ziv Tel Aviv University Nur Lan École Normale Supérieure Emmanuel Chemla École Normale Supérieure Roni Katzir Tel Aviv University
Аннотация
@@ 1 Background: LLMs, HLC, and the inference to the best explanation
Being part of empirical science, the study of HLC proceeds through inference to the best explanation: explicit theories are evaluated in light of the observations, and those theories that provide the best explanations of the data are preferred to the other contenders. The theories coming out of generative linguistics over the past 70 years or so are routinely evaluated in this way. These theories are often strongly linguistically biased, in the sense that they make it relatively easy to represent and learn knowledge of the kind found across languages and hard or impossible to represent or learn many other patterns. Let us use $H_{1}$ to refer to some concrete instantiation of the generative approach. We will return to the biases embodied by the architecture and representations of $H_{1}$ in the following sections.
More recently, Piantadosi (2023) has suggested that current LLMs should be treated as explicit theories of HLC that should be taken seriously as part of this evaluation. Following Fox and Katzir 2024 we refer to this new approach to HLC as the LLM Theory, and we discuss it in Section 1.1, through a concrete instantiation we will call $H_{2}$. Perhaps in light of the glaring inadequacies of the LLM Theory, some recent literature, including Wilcox et al. 2023 and Mahowald et al. 2024, has stopped short of explicitly endorsing the LLM Theory while still taking the performance of current LLMs to challenge central claims within theoretical linguistics. It is conceivable that this more recent literature is still meant to support the LLM Theory, in which case the objections to that theory, reviewed below, still apply. But it is also possible that at least some of these works aim at a more nuanced view on the role of LLMs in studying HLC, one in which the performance of current models can be taken to challenge generative views and support a theory of HLC $H_{3}$ that is not as strongly linguistically-biased as $H_{1}$ but is also very different from current models themselves, and therefore different from $H_{2}$. Our goal in this paper is to articulate this more nuanced view and to make a preliminary attempt to evaluate it in light of current evidence. We also hope to encourage adherents of this view to make explicit the target $H_{3}$ they have in mind.
The idea, as we understand it, is the following. If one wishes to challenge claims for strong linguistic biases, as in $H_{1}$ mentioned above, and to defend a more linguistically-neutral alternative $H_{3}$, one would need to address the arguments that theoretical linguistics has provided for $H_{1}$ from acquisition and from the typology, among other things. But doing so directly can be difficult. For one thing, $H_{3}$ might make it hard to test acquisition predictions based on developmentally-realistic corpora. It can also be hard to evaluate the ability of $H_{3}$ to predict typological patterns. It is at this point that current LLMs can be helpful. While these models might be quite different from $H_{3}$, they too lack the strong linguistic biases that theoretical linguistics has argued for, so their successful performance on tasks related to learning and typology might lend support to the possibility that $H_{3}$ will succeed on these tasks despite not having the strong linguistic biases that theoretical linguistics has argued for. In this way, current LLMs can serve as proxies for $H_{3}$. We refer to this idea as the Proxy View:
[…]
The Proxy View of LLMs in HLC: current LLMs are not themselves good theories of HLC, but their performance is indicative of the success of some other theory of HLC, presumably more linguistically-neutral than the generative approach, and can inform the inference to the best explanation.
Since the literature that promotes the Proxy View generally does not make its supported theory explicit, our ability to evaluate it at present is limited. We will therefore focus on methodological considerations that pertain to the evaluation of the target theory once it is provided and show how LLMs might bear on this evaluation. We show that, at least at present and with respect to the phenomena considered here, LLMs lend no support to the Proxy View.
1.1 The LLM Theory and its inadequacy
As mentioned, Piantadosi (2023) argues for the LLM Theory of HLC. That is, he suggests that current LLMs, taken as real theories of HLC, are a better explanation of the known observations than anything coming out of the generative tradition. This claim is incorrect, for reasons discussed in detail in the literature and summarized briefly here. Following Katzir 2023 and Fox and Katzir 2024 we structure our brief discussion of the failure of the LLM Theory around three key issues: (a) competence vs. performance, (b) correctness vs. probability, and (c) learning and representations. [^1] See Kodner et al., 2023, Moro et al., 2023, Rawski and Baumont, 2023, and others for further critical discussion of the LLM Theory.
Competence vs. performance. Humans struggle with center embedding even at very modest depths of embedding, as in ‘‘The mouse that the cat that the dog chased bit died’’. A naive account might attribute this to the grammar, part of the competence of speakers: perhaps human grammars rule out center embedding beyond a certain (low) threshold. A much better explanation, going back to Yngve 1960 and Miller and Chomsky 1963, attributes the difficulty with center embedding not to the competence of speakers but rather to their performance, typically by attributing it to limitations of their working memory. Among other things, the account in terms of performance explains why center embedding is hard, as seen in reaction times and various errors, as well as from introspection, rather than being just unacceptable. Unacceptability due to the grammar alone, as in ‘‘John ate apple the’’, does not share this property; it is simply judged to be bad. The account in terms of the distinction between competence and performance also explains why the unacceptability of center embedding is gradient, growing with each level of embedding, and why it can vary with certain factors (including with the use of pronouns instead of referring nouns in the most embedded position). Again, the account in terms of competence alone has no explanation for these facts.
Differently from common generative proposals, the design and behaviour of current LLMs does not suggest a meaningful distinction between competence and performance of the kind discovered in humans: there is no reason to think that more (or less) time and working memory would make them change their behaviour on center embedding or agreement. This leaves current LLMs with no good explanation for the increasing difficulty of sentences with increasing levels of center embedding or for the other observations that have been taken to support the distinction between competence and performance.
Correctness vs. likelihood. Humans recognize some sentences as correct but unlikely (one example again is multiple center embedding as in the example above), and they recognize some other sentences as incorrect but likely (for example, so-called agreement attraction, as in the incorrect ‘‘The keys to the cabinet is on the table’’, where is appears in the singular despite a plural subject keys, presumably because of the linear proximity of the singular cabinet). The significance of the distinction between correctness and likelihood for linguistics was pointed out by Chomsky (1957) and discussed further by Berwick (2018), Sprouse et al. (2018), and others.
LLMs provide straightforward information regarding their representation of likelihood: given a prefix, they provide a probability distribution over the next token. But they do not, at present, provide information about correctness as distinct from likelihood, and it remains unclear whether they even have suitable representations to support this distinction. [^2] A potentially important area of research concerns understanding the inner workings of LLMs, particularly through the study of mechanistic interpretation. See Lakretz et al. 2019, Ravfogel et al. 2021, Rai et al. 2024 and López-Otal et al. 2025, among others. Such investigations have the potential of making the inner workings of LLMs less opaque, and it is conceivable that they will eventually uncover a notion of correctness within LLMs.
We note that the literature often compares the likelihood assessments of LLMs to acceptability judgments of speakers. Such, for example, is the methodology of works that evaluate the preferences of LLMs within minimal pairs, including Linzen et al. 2016, Bernardy and Lappin 2017, Chowdhury and Zamparelli 2018, Gulordava et al. 2018, Kuncoro et al. 2018, Marvin and Linzen 2018, Wilcox et al. 2018, 2019, 2023, Bhattacharya and van Schijndel 2020, Chaves 2020, Warstadt et al. 2020, Huebner et al. 2021, Ozaki et al. 2022, Yedetore et al. 2023, and Lan et al. 2024b. This methodology can be informative for particular purposes and under particular conditions, and in Section 2 we adopt it ourselves, but this should not be taken to imply that the relevant likelihood assessments correspond to correctness. As mentioned, the two notions are distinct. [^3] A different method that might appear to offer insight into a notion of correctness or grammaticality in LLMs is prompting, often through interfaces such as ChatGPT. This roughly corresponds to comparing the distribution over ‘yes’ and ‘no’ given a prefix such as “Is the following grammatically correct?” + <target-sentence>. (The precise evaluation depends on the particular prompt that is used and can vary further depending on the specifics of the relevant chat interface.) This is often a convenient way to anecdotally illustrate a general point, by making an LLM appear to engage in reporting its acceptability assessment of the target sentence. While some work, including Dentella et al. 2023 and Qiu et al. 2025, uses prompting on a larger scale, aiming at making general points rather than just illustration, we find this use hard to interpret. In any event, we know of no particular reason to think that prompting provides access to any internal notion of correctness within the LLM (if such a notion even exists).
Learning and representations. The inductive leaps that humans make, sometimes based on very little data, suggest specific representations and biases. This is the essence of so-called arguments from the poverty of the stimulus (APSs). A particularly well-known APS, originally outlined by Chomsky (1971) and discussed in detail in much later work (see Berwick et al. 2011 for a more recent assessment), concerns hierarchical structure, often stated in terms of constituency. [^4] As a reviewer notes, this APS concerns several other syntactic notions in addition to hierarchical structure. For presentational purposes we focus the current discussion on hierarchical structure. This is often illustrated with the formation of yes-no questions, as in (1.1). The generalization that English-learning children reach is that such questions are formed by fronting the auxiliary that is structurally highest, as in (1.1). They do not seem to seriously consider an alternative generalization that says that the fronted auxiliary is the one that is first linearly (which would yield the incorrect (1.1)), even though the data that children are typically exposed to has been argued to be unhelpful in making the choice.
@@ 1.1 The LLM Theory and its inadequacy
Generative theories of HLC typically derive this inductive leap by assuming that children are born with a combination of representations and learning biases that favour dependencies that rely on hierarchical structure over dependencies that rely on linear order (in some approaches by preventing the latter from being represented in the first place). Current LLMs, on the other hand, seem to provide no basis for this inductive leap. In fact, recent work by Lan et al. (2022, 2024a) casts doubt on the ability of networks trained by current methods to even acquire the basic hierarchical notions, let alone to explain the preference for dependencies that rely on these notions over dependencies that rely on linear order. Not surprisingly, then, Yedetore et al. (2023) find evidence suggesting that at least some neural networks fail to acquire the correct hierarchically-based generalization from their training data. We will return to APSs in Section 2, where we will consider the possible role of LLMs in reasoning about the inductive leaps predicted by non-LLM theories of HLC.
@@
Evidence about the representations and learning biases of humans also comes from the highly skewed distribution of linguistic patterns cross-linguistically. We can again consider hierarchical structure, following Fox and Katzir 2024, though we emphasize that a very wide range of cross-linguistic patterns have been discussed in the linguistic literature. We just mentioned that biases that refer to hierarchical structure are part of the best explanation of the competence of English speakers, noting evidence from the formation of yes-no questions. Notions of hierarchical structure are central to the explanation of many other aspects of English, including center-embedding and agreement mentioned earlier, wh-movement discussed in Section 2 below, the computation of meaning, and more. But even if one ignores the APSs for hierarchical structure, it seems clear that hierarchical structure is not just an accident of actual English: it is part of the best explanation of almost all languages that have been studied closely, which suggests that humans are born with representations and learning biases that favour hierarchical structure. Again, the representations and learning biases of current LLMs might not even be able to support the acquisition of the relevant notions of hierarchical structure, let alone explain the cross-linguistic prevalence of these notions. We will return to cross-linguistic evidence in Section 3, where we will consider the possible role of LLMs in reasoning about the typological patterns predicted by non-LLM theories of HLC.
@@ 1.2 LLMs as proxies for good learners
But if the LLM Theory is discredited, does this mean that current LLMs are irrelevant to the scientific study of HLC? As mentioned above, the Proxy View explains why the answer is no: in principle current LLMs can serve as proxies for theories of a third kind, theories that involve relatively linguistically-neutral representations and learning (and are different in this regard from typical generative theories) but that are not vulnerable to the most immediate arguments against the LLM Theory. Table 1 schematically compares the three kinds of theories of HLC that we discuss here: a concrete generative theory, $H_{1}$; a concrete instantiation of the LLM Theory, $H_{2}$; and a concrete, relatively linguistically-neutral, $H_{3}$, possibly very different from current LLMs, for which these LLMs might serve as proxies. The Proxy View, as stated in (1) above, uses current LLMs as tools for evaluating $H_{3}$, in comparison with $H_{1}$ and as part of the inference to the best explanation in light of the existing observations and arguments.
@@
At this point we should mention again an immediate obstacle to evaluating the Proxy View: while recent literature, including Wilcox et al. 2023 and Mahowald et al. 2024, suggests that the performance of LLMs supports linguistically-neutral theories of HLC, this literature does not provide much by way of specifics. Until such specifics of $H_{3}$ are provided, and until $H_{3}$ is shown to be a realistic candidate for the best explanation in light of the existing literature, expressions of hope about $H_{3}$ remain just that.
Still, we can outline some methodological considerations regarding the possible role of current LLMs as proxies for $H_{3}$ once the specifics of $H_{3}$ are made sufficiently explicit. Here we consider two such uses of LLMs as proxies for $H_{3}$. Following Lan et al. 2024b we use LLMs as tools for reasoning about whether $H_{3}$ predicts a given inductive leap based on a given amount of linguistic exposure. And following Kallini et al. 2024 we use LLMs as tools for reasoning about whether $H_{3}$ derives a learning asymmetry that might be part of an explanation of a typological asymmetry.
@@
Explanation for learning asymmetries $\checkmark$ $\times$ $\checkmark$
@@
Таблица 1: Comparison of the three theories outlined above on key issues, following Fox and Katzir 2024.